# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v1

@@ 5 Experimental Results
5.1 Zero-short performance on ImageNet
Table 5 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data. The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column). Models are pre-trained for 30 epochs (pre-train column) and then fine tuned for one epoch (fine-tune column) with the full, unmasked data. Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
We see that CLIPF outperforms the other CLIPA matching approaches in nearly all cases across text masking ratios (input sizes) as reflected by the bolded accuracy scores. In general, random, syntax, and block strategies exhibit similar performance; however, truncation performs worse. Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance. As mentioned in Section 1, we observe that the relative performance of different text masking approaches depends on the number of training epochs. In [12], CLIPA only pre-trained the models for 6.4 epochs. We discuss the learning curve in Fig.1 further in Section 6.2.
@@ 5.1 Zero-short performance on ImageNet
Next we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens). On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios. On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens).
Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens. Here, we focus on the results of the fine-tuned model. Moving from 16 to 8 input tokens benefits most approaches. Recall that the reduction of the number of input tokens allows for an increase of the batch size. Moving from 8 to 6 to 4 tokens, we see that CLIPF maintains its performance surprisingly well. As the masking ratio becomes higher, the margin by which CLIPF outperforms other masking strategies increases. In short, CLIPF is better positioned to deliver benefit as the number of input tokens decreases, reflecting the strength of frequency information in identifying important words to retain during masking.
5.2 Zero-shot performance on different data sets
@@
Following the methodology outlined in CLIP and FLIP [3, 8], we evaluated various masking strategies across data sets to identify their robustness. Table 6 presents results for zero-shot classification for six commonly used data sets. The performance of these strategies on multiple datasets is similar to the behavior observed on the ImageNet-1K dataset. Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases. Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets: ImageNet-A [30] by 1.28%, ImageNet-O [31] by 0.80%, ImageNet-R[32] by 4.9% , ImageNet-Sketch [33] by 3.14%, ImageNet-V2 [34] by 3.38%, and ObjectNet [35] by 0.44%.
Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases. In other words, we observe that the gap in performance widen between CLIP and the other strategies as the masking ratio increases.
We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases. However, despite being the strongest text-matching strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP. In only a handful of cases does it outperform the original CLIP. However, it is important to recall that training with full images, rather than using image masking, can also be combined with text masking. We did not test this setting here for computational reasons, but depending on the computational resources in an application setting it could be relevant.
Table 7: Zero-shot Image-Text Retrieval. We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on COCO and Flickr30k datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@ 5.3 Zero-shot Image-Text Retrieval
Table 7 shows image/text retrieval results on Flickr30k [36] and COCO [37]. These results demonstrate the ability of text masking to improve the performance of a VLM. The performance improvement is particularly notable in comparison to FLIP, which like our CLIPA and CLIPF uses a 75% image-masking ration. When using 16 text tokens, all strategies except the truncation method outperform FLIP which is pre-trained on the full text. The performance is also quite strong in comparison to CLIP, which uses no masking. When using 16 text tokens, text masking strategies, except for truncation, are able to approach and in a few cases beat the accuracy of CLIP.
@@
In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies. The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets, but CLIPF does not consistently outperform the other masking strategies. The performance of the syntax strategy is similar to that of FLIP but is lower than that seen with the random, block, and CLIPF. The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used. Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
6 Analysis and Further Experiments
6.1 Word Distribution Analysis
Figure 2: The figure illustrates the distribution of top-50 words in the text before and after applying various text masking strategies, we set the text length after text masking to 6. The x-axis represents the word index which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$. We remove special characters from the vocabulary.
In this section, we provide an analysis of how different text masking strategies affect the distribution of words By looking at the differences in distribution we can gain insight into why some text-masking strategies work well, and some less well, and, in particular, why CLIPF delivers the best performance.
Fig. 2 shows the distribution of words in the text before and after applying various text masking strategies. We see that in general there is quite a large difference between the frequency of the highest-frequency words and the frequency of words with lower frequency. Such a long-tail distribution is characteristic of words occurring in text, which is known to be Zipf distributed. However, we also observe differences between the different text masking approaches.
The truncation strategy leads to frequent words being more frequent than they are with the other strategies. This effect might seem surprising at first, but can be explained if we consider that texts typically start with the same words, e.g., “A person is…” We conjecture that truncation does not perform well as a text-masking strategy because the VLM overfits to these words.
The CLIPA random and block strategies exhibit nearly identical distributions, since each word has a 50% probability of being masked in both cases. They lie somewhat below the CLIPA truncation line, and we describe them as being more “balanced” not in the sense that the distributions are flat, but rather that the differences between the frequencies of very high-frequency words and the of the rest of the words is less extreme. Even more balanced is the CLIPF distribution, which lies below CLIPA random and block.
The distribution of the CLIPA syntax strategy is quite different. Because this strategy retains more nouns, the articles “the” and “a” or less frequent, but the nouns create spikes. The rationale for prioritizing nouns is convincing. Nouns are closely related to the depicted content of images and for this reason are considered to be more helpful. Indeed, the helpfulness of nouns for the VLM is reflected in the fact that Syntax masking learns faster than other text masking strategies. In Fig. 1 we saw that syntax masking outperforms the other masking approaches during the early epochs and CLIPA [12] found that syntax masking was the best approach, while testing on a limited number of epochs. However, we believe that retaining nouns can have two adverse effects. First, the model can overfit to certain high frequency nouns. In Fig. 2, we see that CLIPA syntax retains more occurrences of “person,” “illustration,” and “stock” than the other strategies. Second, retaining nouns means necessarily dropping words from other part-of-speech (POS) categories.
To dive deeper into the importance of the distribution of words over POS categories, we provide statistics for each of the different text masking strategies in Table 8. Consulting the column ‘Other’ we see and interesting pattern. CLIPA syntax masking changes the proportion of words in the “Other” category radically with respect to the data before masking, but CLIPA truncation, random and block do not. CLIPF appears to provide a happy medium. CLIPF retains more nouns than CLIPA truncation, block, and syntax, but not so many that the “Other” category becomes poorly represented. We believe that it is this balance that this balance is an important factor in CLIPF’s strong performance.
Masking NN (%) JJ (%) VB (%) OTHER (%)
Before masking 50.30% 4.98% 5.18% 39.55%
Truncation 51.72% 5.17% 5.91% 37.20%
Random 51.21% 4.82% 5.04% 38.93%
Block 49.37% 4.74% 5.24% 40.66%
Syntax 88.53% 2.69% 2.92% 5.87%
CLIPF 60.43% 5.06% 6.20% 28.24%
6.2 Learning Curve
Next, we return to consideration of the learning curve that compares CLIPF with the other text-masking strategies, which is shown in Fig. 1. As already mentioned, syntax is a strong performer in the earlier epochs, until epoch 20 remains the best of the CLIPA approaches. We conjecture that CLIPA block and CLIPA random pull ahead of syntax because they retain “Other” words (i.e., words that are not nouns, adjectives or verbs) and do not overprioritize nouns.
The importance of “Other” words is surprising because the category includes many words, which like articles, are said to lack semantic content and have no direct connection to images. In fact, in many NLP applications, articles and other “stop words” are simply removed. We believe that the transformer that CLIP uses to embed text is actually able to exploit the context of content words, particularly words in “Other” category, given enough epochs. CLIPA block masking retains this context a little better than random masking. Surprisingly, the frequency masking of CLIPF also retains this context. Under CLIPA syntax masking a phrase like “light in the house” would always be reduced to “light house”, but with CLIPF frequency masking the reduction would not happen in every case. In short, CLIPF achieves a productive balance in POS categories.
6.3 Comparison with SW-CLIP
Table 9: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k.
Models Datasets Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
SW-CLIP CC3M 49 6 14.9 16.8
CLIPF 14.4 18.2
SW-CLIP CC12M 49 8 35.9 38.5
CLIPF 36.6 39.3
model Image tokens Text tokens Threshold ViT-B/16
pre-train fine-tune
CLIPF 49 8 1e-5 35.6 38.4
1e-6 36.6 39.3
1e-7 36.1 38.6
6.4 Threshold Analysis
7 Conclusion and Outlook
In this paper, we have introduced CLIPF an approach for text masking based on word frequency that improves the performance of CLIP. Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process. Unlike truncation, random, and block text masking strategies, which treat all words the same, CLIPF assigns different probabilities based on word frequency.
In the zero-shot transfer evaluation, CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and the other text masking strategies. Finally, we analyze the word distribution before and after applying text masking to provide insight into why CLIPF is effective. We establish that CLIPF balances retention of nouns, similar to syntax masking, with retention of other parts of speech, similar to block and random masking, effectively providing the best of both worlds.
Future should investigate two directions. First, CLIPF should be tested on larger data sets, that are beyond the reach of our own computational resources. The fact that CLIPF is effective both for CC3M as well as for the larger CC12M data set suggests that CLIPF scales well. We believe that CLIP could be adapted to any larger-scale image-text dataset; however, this should be shown empirically. We see no intrinsic reason why the advantages of CLIPF should not extend to large data sets, in which the frequency of the most frequent words can be expected to be even larger than for small datasets. Second, now that we have gained insights into the distributional properties of text data that appear important for successful text masking, future work can attempt to optimize masking directly to achieve these properties. In the meantime, frequency is what you need.
Appendix A APPENDIX
Appendix B Text Masking Analysis
B.1 Text Masking Cases
Table  11 presents the potential text resulting from the text masking technique. Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text. Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence. As shown in Figure 2, truncation retains more occurrences of “the” and “a” than other text masking strategies. In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency. Syntax masking retains the nouns in the text and remains the same in each epoch. However, this strategy may cause the model to overfit on the frequency of noun words. In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency. This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words. Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others. This helps the model focus on the most helpful aspects of the content. Table 12 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
Table 11: The example provided showcases from the CC12M dataset before and after text masking is applied, reducing the text from 32 tokens to 8 tokens. The first row displays the original text, while the subsequent rows show potential outcomes after applying text masking. The value of $t$ of formular 2 is set to $10^{-6}$.
Text Walk of the happy young couple and Siberian dog. The handsome man is hugging the smiling red head girl
Truncation walk of the happy young couple
Random the happy and the is the
Block couple and siberian dog . the
Syntax walk dog man smiling head girl
CLIPF siberian handsome man hugging smiling red
Word Probability
walk 0.926171
of 0.992838
the 0.995064
happy 0.951695
young 0.957311
couple 0.941174
and 0.991695
siberian 0.75092
dog 0.960531
. 0.993678
B.2 Word Distribution Analysis for CLIPF
Figure 3: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
B.3 Word Distribution of CLIPF and SW-CLIP
We also compare the word distributions between CLIPF and SW-CLIP. As shown in Figure 4, SW-CLIP exhibits a flatter word distribution than CLIPF.
B.4 Word Category Distribution Across Different Text Masking Strategies
As shown in Table 13, we provide the word type counts of Table 8. After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar. However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies. Additionally, SW-CLIP masks a high percentage of other type words, which may impact the model’s zero-shot ability.
Masking NN Count JJ Count VB Count OTHER Count NN (%) JJ (%) VB (%) OTHER (%) Total
Figure  5 shows the word cloud before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying frequency word masking, the word cloud in Table 5(b) shows a more balanced distribution of word frequencies compared to before the masking. The word “the" appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text. As depicted in Table 5(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
Figure 5 shows word clouds before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying word frequency masking, the word cloud in Figure 5(b) displays a more balanced distribution of word frequencies compared to before masking. The word “the” appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text. As depicted in Figure 5(f), syntax masking tends to retain more nouns such as “person”, “stock” and “illustration” resulting in a different pattern from the others.
Refer to caption Refer to caption
(a) Full Text (b) Frequency
Refer to caption Refer to caption
(c) Truncation (d) Random
Refer to caption Refer to caption
(e) Block (f) Syntax
Figure 5: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words. The font size of the word represents the relative word frequency for each figure. The dataset used is CC12M and the value of $t$ of formular 2 is set to $10^{-6}$.
B.5 Bi-gram word Distribution
As shown in Figure 6, we also calculated the bigram distribution of the text after applying different text masking strategies. Truncation and block masking better retain information about the order of the words in the text because the words remain contiguous after masking. Random and CLIPF masking also preserves the text order better than syntax masking. Syntax masking retains more continuous bigrams—such as “stock illustration,” “vector illustration,” and “stock photography”—than other text masking strategies. These bigrams might be domain specific and not contribute to the transferrability of the VLM to other data sets.
Figure 6: The figure illustrates the distribution of top-50 bi-grams in the text with different text masking strategies.
B.6 Impact of Symbolic Characters
We note the high frequency of symbolic characters appearing in the dataset, these symbolic characters are not directly related to the depicted content of the images. Since CLIPF tends to mask these characters with high probability, we removed them during the pre-training of the model to evaluate their significance. As shown in Table 14, removing symbolic characters and stop words does not affect the model’s performance.
Table 14: We remove symbolic characters and stop words during the model’s pre-training to evaluate their zero-shot classification performance on ImageNet-1K. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC3M [23] for 30 epochs with image masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens Remove ViT-B/16
pre-train fine-tune
FLIP 49 32 14.0 14.2
FLIP 49 32 13.9 13.7
B.7 Word and Token Analysis
Since CLIP encodes text using byte-pair encoding (BPE), some words are encoded into more than one token. Therefore, in this experiment, we calculate the masking probability based on tokens rather than words. As shown in Table 15, we compare the performance of both approaches.
Table 15: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification.
Models Masking Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
CLIPF frequency-token 49 8 36.0 38.0
frequency-word 49 8 36.6 39.3
Table 16: Zero-shot accuracy on more classification datasets. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets.
@@ B.7 Word and Token Analysis
Appendix C More results

### v2

@@ 6 Experimental Results
In this section, we carry out a systematic investigation of the existing text masking approaches from CLIPA [12] and our CLIPF approach. Our results support our observation that syntax and truncation masking are not productive approaches to masking and highlight the benefits of masking that directly leverages frequency information, i.e., our CLIPF. Our experiments use 30 epochs, following Open_CLIP [9]. We note that VLM training work has shown performance improvements up to 100 epochs [5, 11] and we would expect that with more epochs the performance would also improve. However, the improvement beyond 30 is expected to be small, predictable and not necessarily important in an operational setting where conserving compute is of concern.
6.1 Zero-shot Performance on ImageNet-1k
Table 4 presents results for zero-shot classification on ImageNet-1K for models trained on both the CC3M and CC12M data. The experiments compare different masking approaches and different masking ratios, indicated in the table with the number of input text tokens (Text Tokens column). Models are pre-trained for 30 epochs (pre-train column) and then fine-tuned for one epoch (fine-tune column) with the full, unmasked data. Note that since the average text length in CC3M is less than 16 tokens, the benefits of fine-tuning a model pre-trained with 16 text tokens are limited.
We see that CLIPF outperforms the other CLIPA matching approaches in nearly all cases across text masking ratios (input sizes) as reflected by the bolded accuracy scores. In general, random, syntax, and block strategies exhibit similar performance; however, truncation performs worse. Contrary to the conclusion of the original CLIPA work [12], syntax masking does not yield the best performance.
@@ 6.1 Zero-shot Performance on ImageNet-1k
Next, we consider the comparison with FLIP, which is the same as CLIP except that it uses a 75% image masking ratio (49 image tokens). On CC3M, fine-tuned CLIPF is competitive with or outperforms fine-tuned FLIP at all masking ratios. On CC12M, fine-tuned CLIPF outperforms fine-tuned FLIP, except for the highest masking ratio (87.5% corresponding to 4 input tokens). Additionally, as shown in the last row of Table 4, if we double the number of image tokens, then CLIPF substantially beats the original CLIP (41.0% vs 36.6%). However, it also notes that more resources are required (Tokens processed from 25% to 46%).
Finally, we consider the pattern of the text masking approaches as the masking ratio increases, i.e., as we decrease the input size from 16 to 4 tokens. Here, we focus on the results of the fine-tuned model. Moving from 16 to 8 input tokens benefits most approaches. Recall that the reduction of the number of input tokens allows for an increase in the batch size. As we reduce the number of tokens from 8 to 6 to 4, syntax masking and truncation do not perform as well as the other text masking strategies. We see that compared to existing text matching approaches, CLIPF is better positioned to deliver benefits as the number of input tokens decreases.
6.2 Zero-shot Robustness Evaluation
@@
Following the evaluation of CLIP and FLIP [3, 8], we test the masking strategies across data sets. Table 6 presents results for zero-shot classification for six commonly used data sets. The performance of these strategies on multiple datasets is similar to the behavior observed on the ImageNet-1K dataset. Across different text masking ratios and across different data sets, CLIPF outperforms the other masking strategies in nearly all cases. Reducing the text tokens in CLIPF from 16 to 8 results in substantial performance improvements compared to the other text masking approaches across almost all datasets: ImageNet-A [31] by 1.28%, ImageNet-O [32] by 0.80%, ImageNet-R[33] by 4.9% , ImageNet-Sketch [34] by 3.14%, ImageNet-V2 [35] by 3.38%, and ObjectNet [36] by 0.44%. Again, we see that the benefits of CLIPF are particularly evident as the masking number of input tokens decreases.
We also observe that CLIPF is able to improve over the FLIP baseline without masking in a majority of cases. However, despite being the strongest text-masking strategy, the results reveal that it is challenging for CLIPF to improve over the original CLIP. In only a handful of cases does it outperform the original CLIP, but might do so with more image tokens as we saw in 6.1.
Table 7: Zero-shot Image-Text Retrieval. We evaluated CLIP, FLIP, CLIPA, and CLIPF image-text retrieval performance on the COCO and Flickr30k datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@ 6.3 Zero-shot Image-Text Retrieval
Table 7 shows image/text retrieval results on Flickr30k [37] and COCO [38]. In general, we see that text masking approaches are able to improve over FLIP. When using 16 text tokens, all strategies except the truncation method outperform FLIP, which is pre-trained on the full text. Here, we do not see consistent improvement over CLIP, although some cases come lose. Note again that we are testing here only with 49 image tokens and 98 would fare better, with additional computational cost.
@@
In contrast to the previous experiments, with image-text retrieval, we observe that CLIPF does not clearly dominate the other text masking strategies. The CLIPA random, CLIPA block, and CLIPF (frequency) strategies exhibit comparable performance in image and text retrieval across both datasets. The notable exception is the performance of CLIPF in the case that the text masking ratio is very high, i.e., only 4 input text tokens are used. Here, the benefit of CLIPF’s frequency-based approach is clearly evident.
We point out that for Image-Text Retrieval the performance of syntax masking is particularly low in comparison to the other forms of masking. Specifically, when using 8 and 6 text tokens, the performance gap between syntax masking and block, random, and frequency masking is approximately 9% on COCO and 6% on Flickr30k for image-text retrieval. This gap is wider than what was observed with Zero-shot Classification. We conjecture that the bad performance is due to the Image-Text Retrieval task requiring understanding the overall meaning and logic of the text rather than just recognizing nouns or objects. For this reason, the fact that syntax masking retains a high number of nouns, as shown in Section 5, is particularly harmful.
Table 8: Zero-shot accuracy on more classification datasets. Comparison of the zero-shot accuracy performance of CLIP, FLIP, CLIPA, and CLIPF on various datasets. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
@@
6.4 Zero-shot Classification on More Datasets
In Table 8, we present evaluation results on additional datasets. With six text tokens, CLIPF consistently outperforms the original CLIP (no masking), FLIP (image masking with full-text), and other text masking strategies. Aligned with ImageNet findings, reducing text tokens from 16 to 8 significantly boosts performance. Improvements include SUN397 [39] (+6.39%), DTD [40] (+9.31%), OxfordPets [41] (+4.64%), STL10 [42] (+9.31%), EuroSAT [43] (+6.60%), UCF101 [44] (+6.68%). The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction. However, their improvements are more limited.
We point out that the Describable Textures Dataset (DTD)[40] contains 47 categories of texture patches, whose POS is mostly verb and that the OxfordPets dataset[41] also contains many class names that are not nouns. On these data sets the performance of syntax masking is particularly low. Specifically, when using six text tokens, syntax masking underperforms CLIPF by 14.10% on DTD and 21.04% on OxfordPets. This low performance reflects that the VLM has overfitted on the large number of nouns retained by syntax masking.
6.5 Scaling to the larger Dataset.
Finally, we carry out an exploratory experiment to gain insight into different masking strategies with data sets of different sizes. We carry out the largest experiment that is possible given our computational capacity, which was training on the LAION-400M [17] data set, using 4 text tokens, a smaller image size (112x112) and 50% image masking. We trained for 16 epochs. Previous works have shown that both smaller image sizes and image masking are effective for accelerating CLIP pre-training [21, 8, 45]. Results are shown in Fig. 4 for both top-1 accuracy (left) and top-5 accuracy (right). We see from these graphs that using $<$ 10 epochs would not provide a complete picture on how the different masking approaches perform relative to each other. Overall, CLIPF, the frequency based approach dominates. However, syntax is performing better than it did on smaller data sets, possibly because the large amount of data is compensating for the overemphasis of syntax masking on nouns. It is important not to conclude on the basis of this experiment that syntax masking will continue to dominate block masking. Looking at the top-5 performance, we see that these two approaches near each other already in the 16 epochs we are able to compute. In sum, this experiment shows that CLIPF is a good performer as data scales up, but that the number of epochs is important in understanding the strength of different masking strategies.
Figure 4: Zero-shot classification accuracy on ImageNet-1K over training epochs for CLIPF and CLIPA strategies when pre-training on LAION-400M datasets. We pre-trained the model using 25 image tokens and 4 text tokens for 16 epochs.
7 Conclusion and Outlook
This paper has provided a systematic overview of text masking approaches for VLM training, which accelerate training and have the potential to also improve performance. In this paper, we demonstrate that training epochs have a substantial impact on the performance of different text masking strategies. We conclude, contrary to the existing state-of-the-art, CLIPA [12], that syntax masking is not the best performing text masking approach and that truncation masking can be avoided. We carry out an analysis of word frequency distributions that shows that frequency is what you need to think about when text masking the training data of VLMs. Specifically, both syntax and truncation have visible effects on the word frequency distribution, making clear why they do not perform well.
We have introduced CLIPF, an approach for text masking based on word frequency that improves the performance of CLIP. Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process. CLIPF is also interesting because it controls frequency directly and in doing so seems to balance nouns with words in the “Other” POS categories, in contrast to other approaches.
We carry out an extensive comparison among existing text matching strategies (CLIPA variants) and CLIPF that revals the strong performance of CLIPF in a variety of zero-shot classification tasks and when scaling to larger data sets. CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and other text masking strategies. It also delivers strong performance in under highly aggressive masking (i.e., using only 4 tokens per training text). Moving forward, the insights gained from this paper about the importance of frequency can be used to further optimize CLIPF to directly optimize the word distribution allowing for more tuning control when trading off VLM performance and VLM training speech during text masking.
Appendix A Text Masking Analysis
A.1 Text Masking Cases
As illustrated in Fig. 5, there is a clear relationship between word masking probability and word frequency. Frequent words have a higher masking probability compared to infrequent ones. Additionally, a smaller threshold $t$ leads to a smaller difference between the masking probabilities of frequent and infrequent words. Therefore, it was necessary to choose a relatively larger threshold to ensure that both frequent and infrequent words are effectively masked.
Figure 5: The curve of Formula 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of formular 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$,$10^{-8}$.
Table  9 presents the potential text resulting from the text masking technique. Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text. Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence. As shown in Fig. 3, truncation retains more occurrences of “the” and “a” than other text masking strategies. In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency. Syntax masking retains the nouns in the text and remains the same in each epoch. However, this strategy may cause the model to overfit on the frequency of noun words. In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency. This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words. Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others. This helps the model focus on the most helpful aspects of the content.
Table 10 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
Table 9: Example from the CC12M dataset illustrating the effect of various text masking strategies, reducing text length to 6 words. The original caption is in the first row, followed by masked variants. Parameter $t$ in Equation 2 is set to $10^{-6}$.
Method Resulting Text
Original Walk of the happy young couple and Siberian dog. The handsome man is hugging the smiling red head girl
Truncation walk of the happy young couple
Random the happy and the is the
Block couple and siberian dog . the
Syntax walk dog man smiling head girl
CLIPF siberian handsome man hugging smiling red
Word Probability
walk 0.926171
of 0.992838
the 0.995064
happy 0.951695
young 0.957311
couple 0.941174
and 0.991695
siberian 0.75092
dog 0.960531
. 0.993678
A.2 Word Category Distribution Across Different Text Masking Strategies
As shown in Table 11, we provide the word type counts of Table 5. After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar. However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies. Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
Masking NN Count JJ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
Figure  6 shows the word cloud before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying frequency word masking, the word cloud in Table 6(b) shows a more balanced distribution of word frequencies compared to before the masking. The word “the" appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text. As depicted in Table 6(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
Fig. 6 shows word clouds before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying word frequency masking, the word cloud in Fig. 6(b) displays a more balanced distribution of word frequencies compared to before masking. The word “the” appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text. As depicted in Fig. 6(f), syntax masking tends to retain more nouns such as “person”, “stock”, and “illustration”, resulting in a different pattern from the others.
Refer to caption Refer to caption
(a) Before masking (b) Frequency
Refer to caption Refer to caption
(c) Truncation (d) Random
Refer to caption Refer to caption
(e) Block (f) Syntax
Figure 6: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words. The font size of the word represents the relative word frequency for each figure. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
A.3 Word Distribution Analysis for CLIPF
Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
A.4 Word Distribution of CLIPF and SW-CLIP
We also compare the word distributions of CLIPF and SW-CLIP after applying text masking. As illustrated in Fig. 8, SW-CLIP masks more frequent words based on its threshold, preventing it from fully utilizing the input slots of the text encoder. Consequently, SW-CLIP exhibits a flatter word distribution compared to CLIPF.
Appendix B More Results